OpenStudy (anonymous):

the rectangular coordinates are given, find polar coordinates in (r,theta) of each point (-4,4 sqrt3) a.use the equation r^2=x^2+y^2 to find the radius of the point b.use the fact that tantheta=y/x, x is not equal to 0 to find the angle (in radians) for this point b.combine your answer to write polar coordinates of this point

OpenStudy (anonymous):

\[r^2 = (-4)^2 + (4 \sqrt 3)^2 = 16 +48 =64\\ r=8 \]

OpenStudy (anonymous):

\[ \tan (\theta) =-\frac { 4 \sqrt 3} 4= -\sqrt 3\\ \theta=-\frac \pi 3 \]